# Deep Ensemble Shape Calibration: Multi-Field Post-hoc Calibration in Online Advertising

### v1

@@ Abstract.
In the e-commerce advertising scenario, estimating the true probabilities (known as a calibrated estimate) on Click-Through Rate (CTR) and Conversion Rate (CVR) is critical and can directly affect the benefits of the buyer, seller and platform.
Previous research has introduced numerous solutions for addressing the calibration problem. These methods typically involve the training of calibrators using a validation set and subsequently applying these calibrators to correct the original estimated values during online inference. However, what sets e-commerce advertising scenarios is the challenge of multi-field calibration. Multi-field calibration can be subdivided into two distinct sub-problems: value calibration and shape calibration. Value calibration is defined as no over- or under-estimation for each value under concerned fields (such as the average of pCTR should equal to the CTR for value ”women’s shoes” in the field ”category”, in the advertising context, the number of fields can range from dozens to even hundreds). Shape calibration is defined as no over- or under-estimation for each subset of the pCTR within the specified range under condition of concerned fields. In order to achieve shape calibration and value calibration, it is necessary to have a strong data utilization ability. Because the quantity of pCTR specified range for single field-value (such as user ID and item ID) sample is relative small, which makes the calibrator more difficult to train. However the existing methods cannot simultaneously fulfill both value calibration and shape calibration.
To solve these problems, we propose a new method named Deep Ensemble Shape Calibration (DESC)  [^1] The code is at https://github.com/HaoYang0123/DESC. We introduce innovative basis calibration functions, which enhance both function expression capabilities and data utilization by combining these basis calibration functions. A significant advancement lies in the development of an allocator capable of allocating the most suitable shape calibrators to different estimation error distributions within diverse fields and values. Finally, we achieve significant improvements in both public data and industrial data, and on online experiments we get +2.5% of CVR and +4.0% of GMV (Gross Merchandise Volume).
@@ 1. INTRODUCTION
However, recent studies (Bella et al., 2010; Guo et al., 2017a; Korb, 1999; Ovadia et al., 2019) have revealed that numerous established machine learning techniques, particularly deep learning methods extensively applied in the fields of e-commerce advertising, often yield inadequately calibrated probability predictions.
@@
However, what sets apart e-commerce advertising scenarios is the challenge of multi-field calibration, which means that for each pCTR range in each value under each concerned field the pCTR should be calibrated. Multi-field calibration can be subdivided into two distinct sub-problems: value calibration and shape calibration. Value calibration is defined as no over- or under-estimation each value under concerned fields (such as average of pCTR should equal to the CTR for value ”women’s shoes” in the field ”category”, in the e-commerce advertising context, the number of fields can range from dozens to even hundreds). Figure 1 illustrates the significant inconsistency of over- and under-estimation across different field values. Shape calibration is defined as no over- or under-estimation for each subset of the pCTR within the specified range for each value under concerned fields. Take Figure 2 as an example, Field $f$ has two values: Value1 and Value2, and these two shapes are significantly different. In order to achieve shape calibration and value calibration, it is necessary to have a strong data utilization ability. Because the samples of the specified pCTR range for single field-value (such as user ID and item ID) are relative small, which makes the calibrator more difficult to train. However, the existing method cannot simultaneously fulfill both value calibration and shape calibration.
For Parameter methods, non-parameter methods and non-field-aware methods, they involve the training of a single calibration function to address the issue of shape miscalibration. These methods focus on shape calibration but not value calibration.
@@
  • Two public datasets and one industrial dataset are tested to show that our proposed DESC method achieves significantly better performance compared with other methods in both calibration metrics and non-calibration metrics. And we get +2.5% of CVR and +4.0% of GMV on the online experiment.
@@ 2.1. Non-parametric Methods
Non-parametric methods, in contrast, do not rely on any assumptions regarding the distribution of estimates. Examples of non-parametric methods encompass Histogram Binning (HB)(Zadrozny and Elkan, 2001) and Isotonic Regression (IR) (Zadrozny and Elkan, 2002). HB involves sorting the estimated values and then dividing them into bins of equal frequencies or intervals, utilizing the posterior CTR of each bin as a calibration parameter. Building upon this, IR introduces the additional constraint of order preservation. These approach can introduce instability in the bin boundary values.
@@ 2.3.1. Non-field-aware Methods
Non-field-aware methods encompass three notable approaches: Scaling-binning (Kumar et al., 2019), Smooth Isotonic Regression (SIR) (Jiang et al., 2011), and Ensemble Temperature Scaling (ETS) (Zhang et al., 2020). Scaling-binning combines the techniques of Platt Scaling and Histogram Binning. SIR employs linear interpolation based on isotonic regression, while ETS combines multiple temperature scalings for calibration, offering greater expressiveness than temperature scaling. Notably, all of these methods utilize raw predicted scores as input without taking field information into account.
@@ 2.3.2. Field-aware Methods
Field-aware methods integrate supplementary features to formulate calibration functions. NeuralCalib (FAC) (Pan et al., 2020) introduces an auxiliary module by utilizing neural networks and incorporates multiple field modeling. However, it does not address the challenge of shape miscalibration variance among different field values. AdaCalib (Ada) (Wei et al., 2022) learns an isotonic function based on posterior statistics and selects the most appropriate bin number for a single field value. However, AdaCalib does not encompass all fields. MBCT (Huang et al., 2022) employs trees to uncover more effective calibration across fields. However, due to its high computational complexity, calibration trees are unable to handle sparse fields, such as user ID and item ID, and do not adequately address discrepancies in shape calibration.
3. PROBLEM FORMULATION
In the context of online advertising systems, using CTR as an illustration (CVR follows the same principles), we have trained a neural predictor, denoted as $f_{uncalib}$, on a training dataset $\mathcal{D}_{train}$. This dataset includes all field values as inputs ($x$) and click responses ($y$), where $y=0$ represents non-click events and $y=1$ signifies click events. The neural predictor is capable of forecasting the likelihood of a click using the following formula.
@@ 4. METHODS
(3) $\hat{p}_{calib}=\mathbb{S}(x)\cdot\mathbb{V}(x).$
$\mathbb{S}(x)$ and $\mathbb{V}(x)$ refer to the shape calibrated score (the output of Shape Calibrator) and value calibrated score (the output of Value Calibrator), respectively. The input $x$ consists of all $n$ field values and non-calibrated scores $\hat{p}_{uncalib}$. The training loss function is the negative log-likelihood function.
@@ 4.1.1. Shape Pre-Define
Basis calibration function $\mathcal{B}(x)$ satisfies the following conditions: 1. The function is monotonically non-decreasing and continuous in range $(0,1)$.
2. When $x$ approaches $0^{+}$, the limit of $\mathcal{B}(x)$ is 0, and when $x$ approaches $1^{-}$, the limit of $\mathcal{B}(x)$ is 1 (Equation 5).
(5) $\lim_{x\to 0^{+}}\mathcal{B}(x)=0\ and\ \lim_{x\to 1^{-}}\mathcal{B}(x)=1$
In Equation 6, we pre-define $m$ basis functions consisting of $p$ power functions, $l$ logarithmic functions and $s$ scaling functions. These functions exhibit different shape characteristics, and their shapes vary when the hyper-parameters (e.g. the coefficients in these basis calibration functions) are different. Compared to segmented linear functions in traditional calibration methods, including SIR, FAC, Ada, using basis calibration functions doesn’t require data segmentation. In other words, shape learning can utilize all samples for training, significantly enhancing the data utilization.
(6) $\begin{matrix}\par\mathcal{B}(x)=\{\underbrace{\mathcal{B}_{1}^{power},...,\mathcal{B}_{p}^{power}}_{p\ power\ functions},\underbrace{\mathcal{B}_{1}^{log},...,\mathcal{B}_{l}^{log}}_{l\ log\ functions},\underbrace{\mathcal{B}_{1}^{scaling},...,\mathcal{B}_{s}^{scaling}}_{s\ scaling\ functions}\}\end{matrix}$
(7) $\mathcal{B}_{i}^{power}(x;h_{i})=x^{h_{i}}$
(8) $\mathcal{B}_{i}^{log}(x;v_{i})=\frac{log(1+v_{i}\cdot x)}{log(1+v_{i})}$
(9) $\mathcal{B}_{i}^{scaling}(x;a_{i})=\sigma(\sigma^{-1}(x)\cdot a_{i})\ \ ,$
(10) $where\ \sigma(x)=\frac{1}{1+exp(-x)}\ and\ \sigma^{-1}(x)=log(\frac{x}{1-x})$
$h_{i},v_{i}$ and $a_{i}$ (${h_{i},v_{i},a_{i}}\in\mathbf{R}^{+}$) are parameters of basis calibration functions. We pre-define these parameters using equally spaced floats (e.g. $[0.3,0.5,0.7,0.9]$) for each type of basis calibration function.
However, $\mathcal{B}(x)$ can only represent simple shapes, and complex shapes can be composed of simple shapes, as shown in Figure 4. Therefore, we need to combine these basis calibration functions through weighted summation to create shape function $\mathbb{S}_{i}(x)$ capable of representing complex shapes for $i$-th field, as shown in Equation 11, where $\alpha_{i}^{j}$ refers to the attention weight of $j$-th basis function for $i$-th field, which will be explained in next section.
(11) $\mathbb{S}_{i}(x)=\sum_{j=1}^{m}{\alpha_{i}^{j}\cdot\mathcal{B}_{j}(x)},\;\;\mathcal{B}_{j}(x)\in\mathcal{B}(x)$
@@ 4.1.2. Shape Allocation
In shape allocation, we allocate suitable shape functions based on feature values. These features include two parts: pCTR bucket feature and original field features. Introducing pCTR bucket feature allows for better shape allocation within different pCTR intervals. The specific steps are as follows:
@@ 4.2.1. Global Field Value Calibrator
Global Field Value Calibrator encompasses all the necessary information of the fields concerned. These input fields are projected into fixed-size dense representations, which are fed into a neural network (the form of the neural network is not restricted, here we use MLP). Then we get the middle output $hiddenLayer$ and the final output $\mathbb{V}(x)$ for the Global Field Value Calibrator. Finally, by combining $\mathbb{V}(x)$ and $\mathbb{S}(x)$, we obtain the final calibrated output $\hat{p}_{calib}=\mathbb{S}(x)\cdot\mathbb{V}(x)$.
@@ 4.2.2. Global Shape Attention Generator
The Global Field Value Calibrator not only produces the final output but also yields valuable intermediate representation. In this context, we use the intermediate representation $hiddenLayer$ as input of Equation 18, which can effectively learn information from all fields. After passing through a MLP and softmax layer, it produces Global Shape Attention Score $\Psi$ with a size equal to the number of fields $n$. In the Shape Calibration stage, the output produced by each field $Z_{i}$ is multiplied by its corresponding score $\Psi_{i}$ and summed to obtain the final Shape Calibration Score.
@@ 4.3. Online Deployment
The overall framework of the online service using the DESC calibration method in our industrial advertising system is shown in Figure 5. When a real-time user request comes, all candidate items are recalled and predicted with pCTRs by a non-calibration model (ranking model). Then DESC model calibrates the pCTRs considering different field values, including user, item, and context features. In detail, the input $x$ of DESC has two parts: different field values and its non-calibrated score $\hat{p}_{uncalib}$. DESC uses a light neural network to output the calibrated score $\hat{p}_{calib}$, which is then used to sort the final ranks for all candidate items. We have integrated the DESC method as a plugin into the ranking model to reduce maintenance costs while only adding a little more time for online inference.
@@ 5.1. Experimental Setup
To validate the effectiveness of our proposed DESC method, we conduct experiments on two public datasets (AliCCP and CRITEO for CTR prediction tasks) and one industrial dataset (Shopee for CVR prediction task). CRITEO display advertising data 222 https://www.kaggle.com/c/criteo-display-ad-challenge, which consists of 46 million samples. We split the samples into 28 million samples as training set ($\mathcal{D}_{train}$), 9 million samples as validation set ($\mathcal{D}_{dev}$) and other 9 million samples as testing set ($\mathcal{D}_{test}$). AliCCP (Alibaba Click and Conversion Prediction) (Ma et al., 2018) 333 https://tianchi.aliyun.com/datalab/dataSet.html?dataId=408 contains 80 million samples. We split these samples with the proportion 2:1:1 to $\mathcal{D}_{train}$, $\mathcal{D}_{dev}$ and $\mathcal{D}_{test}$. To test DESC method on CVR prediction task, we collect the conversion logs from the Shopee’s online advertising system. It contains 100 million samples, where the first 60 million for $\mathcal{D}_{train}$, the next 20 million for $\mathcal{D}_{dev}$ and the last 20 million for $\mathcal{D}_{test}$.
@@
We use DeepFM (Guo et al., 2017b) to train the non-calibrated models with all fields in $\mathcal{D}_{train}$, and predict the non-calibrated scores in $\mathcal{D}_{dev}$ and $\mathcal{D}_{test}$.

### v2

@@ Abstract.
In the e-commerce advertising scenario, estimating the true probabilities (known as a calibrated estimate) on Click-Through Rate (CTR) and Conversion Rate (CVR) is critical. Previous research has introduced numerous solutions for addressing the calibration problem. These methods typically involve the training of calibrators using a validation set and subsequently applying these calibrators to correct the original estimated values during online inference.
However, what sets e-commerce advertising scenarios is the challenge of multi-field calibration. Multi-field calibration requires achieving calibration in each field. In order to achieve multi-field calibration, it is necessary to have a strong data utilization ability. Because the quantity of pCTR specified range for single field-value (such as user ID and item ID) sample is relatively small, which makes the calibrator more difficult to train. However, existing methods have difficulty effectively addressing these issues.
To solve these problems, we propose a new method named Deep Ensemble Shape Calibration (DESC). In terms of business understanding and interpretability, we decompose multi-field calibration into value calibration and shape calibration. We introduce innovative basis calibration functions, which enhance both function expression capabilities and data utilization by combining these basis calibration functions. A significant advancement lies in the development of an allocator capable of allocating the most suitable calibrators to different estimation error distributions within diverse fields and values. We achieve significant improvements in both public and industrial datasets. In online experiments, we observe a +2.5% increase in CVR and +4.0% in GMV. (Gross Merchandise Volume). Our code is now available at: https://github.com/HaoYang0123/DESC.
@@ 1. INTRODUCTION
Figure 1. Examples to show the shape miscalibration and value miscalibration
However, recent studies (Bella et al., 2010; Guo et al., 2017a; Korb, 1999; Ovadia et al., 2019; Feldman et al., 2023) have revealed that numerous established machine learning techniques, particularly deep learning methods extensively applied in the fields of e-commerce advertising, often yield inadequately calibrated probability predictions.
@@
In the advertising scenarios, there are multiple fields (such as user group and item category), and if predictions of CTR (pCTR) or CVR (pCVR) are not calibrated in certain fields, such as item category, it can lead to a negative impact on the earnings of some advertisers. Simultaneously, in the e-commerce scenario, there is a particularly large number of fields to consider. This necessitates the calibration of each field, referred to as multi-field calibration. To achieve multi-field calibration, it is necessary to have a strong data utilization ability. Because the samples of the specified pCTR range for a single field value (such as user ID and item ID) are relatively small, which makes the calibrator more difficult to train.
In terms of business understanding and interpretability, we decompose multi-field calibration into value calibration and shape calibration. Figure 1 shows shape miscalibration and value miscalibration. Value calibration is defined as no over- or under-estimation each value under concerned fields (such as average of pCTR should equal to the CTR for value ”women’s shoes” in the field ”category”, in the e-commerce advertising context, the number of fields can range from dozens to even hundreds). From the advertising perspective, value calibration ensures that the ECPM (Effective Cost Per Mille) and GMV (Gross Merchandise Value) of different items are not over- or under estimated. Figure 2 illustrates the significant inconsistency of over- and under-estimation across different values. Shape calibration is defined as no over- or under-estimation for each subset of the pCTR within the specified range. Explaining from the advertising perspective, shape calibration ensures that some already popular items are not excessively exposed or suppressed. Take Figure 3 as an example, for the field ”Category” with values ”Food” and ”Book”, the calibrated pCTR/CTR is 1 in any pCTR interval, whereas the distribution (shape) of the uncalibrated pCTR/CTR will be different. However, the existing methods cannot simultaneously fulfill both value calibration and shape calibration.
For parameter methods, non-parameter methods and non-field-aware methods, they involve the training of a single calibration function to address the issue of over- or under-estimation for each subset of the pCTR globally but overlook the biases across different fields.
@@
  • Our proposed DESC method outperforms other methods significantly across both calibration and non-calibration metrics, as demonstrated on two public datasets and one industrial dataset. Additionally, in online experiments, we observe a notable increase of +2.5% in CVR and +4.0% in GMV.
@@ 2.1. Non-parametric Methods
Non-parametric methods do not rely on any assumptions regarding the distribution of estimates. Examples of non-parametric methods encompass Histogram Binning (HB)(Zadrozny and Elkan, 2001) and Isotonic Regression (IR) (Zadrozny and Elkan, 2002). HB involves sorting the estimated values and then dividing them into bins of equal frequencies or intervals. Building upon this, IR introduces the additional constraint of order preservation. These approach can introduce instability in the bin boundary values.
@@ 2.3.1. Non-field-aware Methods
Non-field-aware methods encompass three notable approaches: Scaling-binning (Kumar et al., 2019), Smooth Isotonic Regression (SIR) (Jiang et al., 2011), and Ensemble Temperature Scaling (ETS) (Zhang et al., 2020). Scaling-binning combines the techniques of Platt Scaling and Histogram Binning. SIR employs linear interpolation based on isotonic regression, while ETS combines multiple temperature scalings for calibration. Notably, all of these methods utilize raw predicted scores as input without taking field information into account.
@@ 2.3.2. Field-aware Methods
Field-aware methods integrate supplementary features to formulate calibration functions. NeuralCalib (FAC) (Pan et al., 2020) introduces an auxiliary module but it does not address the challenge of shape miscalibration variance among different field values. AdaCalib (Ada) (Wei et al., 2022) learns an isotonic function based on posterior statistics and selects the most appropriate bin number for a single field value. However, AdaCalib does not encompass all fields. MBCT (Huang et al., 2022) employs trees to uncover more effective calibration across fields. However, calibration trees are unable to handle sparse fields, such as user ID and item ID.
3. CALIBRATION PROBLEM FORMULATION
In online advertising systems, using CTR as an illustration (CVR follows the same principles), we have trained a neural predictor, denoted as $f_{uncalib}$, on a training dataset $\mathcal{D}_{train}$. This dataset includes all field values as inputs ($x$) and click responses ($y$), where $y=0$ represents non-click events and $y=1$ signifies click events. The neural predictor is capable of forecasting the likelihood of a click using the following formula.
@@ 4. METHODS
(3) $\hat{p}_{calib}=\mathbb{S}(x)\cdot\mathbb{V}(x)\ ,\ x=\{Z,\hat{p}_{uncalib}\}$
The input $x$ consists of all $n$ field values $Z$ and non-calibrated scores $\hat{p}_{uncalib}$. $\mathbb{S}(x)$ and $\mathbb{V}(x)$ refer to the shape calibrated score (the output of Shape Calibrator) and value calibrated score (the output of Value Calibrator), respectively. The training loss function is the negative log-likelihood function.
@@ 4.1.1. Shape Pre-Define
For the variable $t$, $t$ is between 0 and 1, basis calibration function $\mathcal{B}(t)$ satisfies the following conditions:
1. The function is monotonically non-decreasing and continuous in range $(0,1)$.
2. When $t$ approaches $0^{+}$, the limit of $\mathcal{B}(t)$ is 0, and when $t$ approaches $1^{-}$, the limit of $\mathcal{B}(t)$ is 1 (Equation 5).
(5) $\lim_{t\to 0^{+}}\mathcal{B}(t)=0\ and\ \lim_{x\to 1^{-}}\mathcal{B}(t)=1$
We pre-define $m$ basis functions consisting of $p$ power functions, $l$ logarithmic functions and $s$ scaling functions (shown in Equation 6). These functions exhibit different shape characteristics, and their shapes vary when the hyper-parameters (e.g. the coefficients in these basis calibration functions) are different. We choose predefined functions over Multi-Layer Perceptron (MLP) because, through data research, we have found that predefined functions can address the issue of over- or underestimation for each subset of pCTR within the specified range (shape calibration). Additionally, predefined functions have a lower parameter count and superior performance.
Compared to segmented linear functions in traditional calibration methods, including SIR, FAC, Ada, using basis calibration functions doesn’t require data segmentation. In other words, shape learning can utilize all samples for training, significantly enhancing the data utilization. Basis calibration functions include, but are not limited to $log$, $exp$ and $scaling$. By analyzing the data, we find that the combination of $log$, $exp$ and $scaling$ functions can satisfactorily fulfill our requirements with ease.
(6) $\begin{matrix}\par\mathcal{B}(t)=\{\underbrace{\mathcal{B}_{1}^{power},...,\mathcal{B}_{p}^{power}}_{p\ power\ functions},\underbrace{\mathcal{B}_{1}^{log},...,\mathcal{B}_{l}^{log}}_{l\ log\ functions},\underbrace{\mathcal{B}_{1}^{scaling},...,\mathcal{B}_{s}^{scaling}}_{s\ scaling\ functions}\}\end{matrix}$
(7) $\mathcal{B}_{i}^{power}(t;h_{i})=x^{h_{i}}$
(8) $\mathcal{B}_{i}^{log}(t;v_{i})=\frac{log(1+v_{i}\cdot t)}{log(1+v_{i})}$
(9) $\mathcal{B}_{i}^{scaling}(t;a_{i})=\sigma(\sigma^{-1}(t)\cdot a_{i})\ \ ,$
(10) $where\ \sigma(t)=\frac{1}{1+exp(-t)}\ and\ \sigma^{-1}(t)=log(\frac{t}{1-t})$
$h_{i},v_{i}$ and $a_{i}$ (${h_{i},v_{i},a_{i}}\in\mathbf{R}^{+}$) are parameters of basis calibration functions. For each type of basic calibration function, we pre-define these parameters using equally spaced floats (e.g., $[0.1,0.3,0.5,0.7]$), which can be set as trainable.
However, $\mathcal{B}(t)$ can only represent simple shapes, and complex shapes can be composed of simple shapes, as shown in Figure 5. Therefore, we need to combine these basis calibration functions through weighted summation to create shape function $\mathbb{S}_{i}(x)$ capable of representing complex shapes for $i$-th field, as shown in Equation 11, where $\alpha_{i}^{j}$ refers to the attention weight of $j$-th basis function for $i$-th field, which will be explained in next section.
(11) $\mathbb{S}_{i}(x)=\sum_{j=1}^{m}{\alpha_{i}^{j}\cdot\mathcal{B}_{j}(t)},\;\;\mathcal{B}_{j}(t)\in\mathcal{B}(t)$
@@ 4.1.2. Shape Allocation
In shape allocation, we allocate suitable shape functions based on feature values. These features include two parts: pCTR bucket feature and original field features. PCTR bucket feature categorizes pCTR into intervals. For example, 0 to 0.001 is bucket 1, and 0.001 to 0.003 is bucket 2. Introducing pCTR bucket feature allows for better shape allocation within different pCTR intervals. The specific steps are as follows:
@@ 4.2.1. Global Field Value Calibrator
Global Field Value Calibrator encompasses all the necessary information of the fields concerned. We train a neural network to fix the field-level miscalibration or biases by utilizing all necessary features(Pan et al., 2020). These input fields are projected into fixed-size dense representations, which are fed into a neural network (the form of the neural network is not restricted, here we use MLP). Then we get the middle output $hiddenLayer$ and the final output $\mathbb{V}(x)$ for the Global Field Value Calibrator. Finally, by combining $\mathbb{V}(x)$ and $\mathbb{S}(x)$, we obtain the final calibrated output $\hat{p}_{calib}=\mathbb{S}(x)\cdot\mathbb{V}(x)$.
@@ 4.2.2. Global Shape Attention Generator
The Global Field Value Calibrator not only produces the final output but also yields valuable intermediate representation. In this context, we use the intermediate representation $hiddenLayer$ as input of Equation 18, which can effectively learn information from all fields. After passing through a MLP and softmax layer, it produces Global Shape Attention $\Psi$ with a size equal to the number of fields $n$. In the Shape Calibration stage, the output produced by each field $Z_{i}$ is multiplied by its corresponding score $\Psi_{i}$ and summed to obtain the final Shape Calibration Score.
@@ 4.3. Online Deployment
The overall framework of the online service using the DESC calibration method in our industrial advertising system is shown in Figure 6. When a real-time user request comes, all candidate items are recalled and predicted with pCTRs by a non-calibration model (ranking model). Then DESC model calibrates the pCTRs considering different field values, including user, item, and context features. In detail, the input $x$ of DESC has two parts: different field values and its non-calibrated score $\hat{p}_{uncalib}$. DESC uses a light neural network to output the calibrated score $\hat{p}_{calib}$, which is then used to sort the final ranks for all candidate items. We have integrated the DESC method as a plugin into the ranking model to reduce maintenance costs while only adding a little more time (about 2ms for one user’s request) for online inference.
@@ 5.1. Experimental Setup
To validate the effectiveness of our proposed DESC method, we conduct experiments on two public datasets (AliCCP and CRITEO for CTR prediction tasks) and one industrial dataset (Shopee for CVR prediction task). CRITEO display advertising data 111 https://www.kaggle.com/c/criteo-display-ad-challenge, which consists of 46 million samples. We split the samples into 28 million samples as training set ($\mathcal{D}_{train}$), 9 million samples as validation set ($\mathcal{D}_{dev}$) and other 9 million samples as testing set ($\mathcal{D}_{test}$). AliCCP (Alibaba Click and Conversion Prediction) (Ma et al., 2018) 222 https://tianchi.aliyun.com/datalab/dataSet.html?dataId=408 contains 80 million samples. We split these samples with the proportion 2:1:1 to $\mathcal{D}_{train}$, $\mathcal{D}_{dev}$ and $\mathcal{D}_{test}$. To test DESC method on CVR prediction task, we collect the conversion logs from the Shopee’s online advertising system. It contains 100 million samples, where the first 60 million for $\mathcal{D}_{train}$, the next 20 million for $\mathcal{D}_{dev}$ and the last 20 million for $\mathcal{D}_{test}$.
@@
We use DeepFM (Guo et al., 2017b) to train the non-calibrated models with all fields in $\mathcal{D}_{train}$, and predict the non-calibrated scores in $\mathcal{D}_{dev}$ and $\mathcal{D}_{test}$. DeepFM consists of both the fully connected part and the FM part. We utilized the open-source DeepFM ”from deepctr.models import DeepFM” and incorporated all features into DeepFM.